The length of a rectangle is four less than three times another. If the perimeter is 26 inches, find the length and width.
Answer by mananth(16946)

let width be x
length = 3x-4
perimeter = 26
perimeter = 2(L+W)
26=2*(x+3x-4)
26=2(4x-4)
26=8x-8
8x=26+8
8x=34
x= 34/8
x= 4.25 inches
length = 3x-4
=12.75-4
=8.75inches
..
m.ananth@hotmail.ca